# Perturbation Theory and the Quantum Rabi-model

### v1

@@ Abstract.
In the first part of the paper we study a perturbative model of the Rabi system of Quantum Optics. We therefore are able to describe, through Rellich’s theory, an analytic expansion of finite families, but of any fixed length, of eigenvalues. In particular, we prove that for finite families of eigenvalues the Braak conjecture holds. In the second part we study the asymptotics of the Weyl spectral counting function of a class of systems that generalize the Quantum Rabi Model to an $N$-level atom ($N\geq 3$) with N1N-1 cavity modes of the electromagnetic field.
@@ 1. Introduction
Being a system on $\mathbb{R}^{n}$ with polynomial coefficients, based on the Quantum Harmonic Oscillator, it fits in the framework of Non-Commutative Harmonic Oscillators (NCHO) introduced by Wakayama and Parmeggiani in [17, 18] and as extended in Parmeggiani [15, 16], and more recently in Malagutti and Parmeggiani [12] (and [13]). However, a difficulty in the geometric analysis of the Rabi system is that it is not a semiregular globally elliptic system in the sense of [12], in that, although the symbol is principally scalar, the symbol of the first order part (the semiprincipal symbol) cannot be smoothly diagonalized in the phase space $\mathbb{R}^{n}\times\mathbb{R}^{n}$.
@@
The paper is divided into two parts, both of which are perturbative in nature. In the first part we study a perturbation of the classical $1$-d QRM, which can be dealt with by the Rellich perturbation theory as in [21] (we chose that approach in place of Kato’s one [9] because more explicit). In this part we destroy the homogeneity of the system, which allows us to bring the system into an isometric system to which we may apply perturbation theory. Since all the eigenvalues of the unperturbed system (a scalar quantum harmonic oscillator acting on $L^{2}(\mathbb{R};\mathbb{C}^{2})$) are degenerate, we apply Rellich’s approach by assuming some first-order (in the perturbation parameter) nondegeneracy, described by the way the interaction parameter is approximated by zeros of Legendre polynomials. That allows to control finite initial segments of the spectrum of the system, for a sufficiently small perturbation parameter, and to prove that in this context Braak’s Conjecture (see Section 5) holds, at least for arbitrarily long but finite parts of the spectrum. We also deal with ”degenerate” (in the sense of Rellich) situations, but in that case we use the method of quasimodes (see Lazutkin [10]). In the end, Braak’s conjecture keeps holding. In a sense, our result may be thought of as a complement of Rudnik’s result [22]. For generalizations of the Braak’s Conjecture to asymmetric quantum Rabi models, see Braak, Nguyen, Reyes-Bustos and Wakayama [3].
@@
In the second part of the paper we describe in the first place some possible generalizations of the QRM to higher dimensions for an interaction with atoms having several energy levels, where the space dimension $n$ is related with the number $N$ of atomic levels by the equation n=N1.n=N-1. In order to approach the problem of the asymptotics of the Weyl spectral counting function, which is related to the non-diagonalizability in the symbol class of the first order part of the system, we use a perturbation method that brings the perturbed symbol into the SMGES class, for which we may exploit a pseudodifferential diagonalization of the system and the results of [12]. Variational considerations then allow us to give a two-term asymptotics of the Weyl function. The result is not known in the higher dimensional and energy level number case.
@@
In Section 2 we describe the Quantum Rabi Model from a physical perspective, highlighting its connection with the Jaynes-Cummings model by Rotating Wave Approximation and its importance as a foundational model in Quantum Optics. In Section 3 we recall the setup, that contains also a review of the global pseudodifferential calculus used especially in the final Section 7. In Section 4 we give the perturbative result, under a non-degeneration condition, on finite initial segments of the spectrum of the QRM. In Section 5 we recall Braak’s conjecture and show its validity in our setting. We next deal with the degenerate case by the use of quasi-modes, which is the content of Section 6. In the final Section 7 we introduce the generalizations of the QRM and give a perturbative proof of a two-term asymptotics of the Weyl spectral counting function.
@@ 2. Physics of the Quantum Rabi Model
The Quantum Rabi Model (QRM) describes the fundamental interaction between a two-level quantum system (a qubit, atom, or spin-½) and a single bosonic mode (a quantized harmonic oscillator, typically a cavity or vibrational mode). It plays a central role in Quantum Optics, cavity and circuit QED, trapped ions, and modern quantum technologies, as it captures the minimal setting in which light-matter interaction is in full treated quantum mechanically.
@@ 2.2. Symmetry and integrability
which commutes with the Hamiltonian. This symmetry splits the Hilbert space into even and odd parity subspaces and is essential for understanding the structure of the spectrum. Braak [1] showed that the QRM is integrable in a generalized sense and that its spectrum can be obtained from the zeros of transcendental functions having no simple closed-form, but built from infinite series and special functions.
@@ 2.3.1. Weak coupling: Jaynes-Cummings regime
For $g\ll\omega,\Omega$ and near resonance $\omega\approx\Omega$, the rapidly oscillating counter-rotating terms can be neglected under the rotating-wave approximation (RWA). The Hamiltonian reduces to
@@ 2.3.2. Ultrastrong coupling
When $g/\omega\sim 0.1$, the counter-rotating terms become relevant and the RWA breaks down. This is the ultrastrong coupling regime, experimentally realized in superconducting circuit QED (see the paper by Niemczyk [14]). The ground state contains virtual photons and the system exhibits phenomena such as the Bloch–Siegert shift.
@@ 3. The setup
We work in the phase-space $T^{\ast}\mathbb{R}^{n}=\mathbb{R}^{n}\times\mathbb{R}^{n}=\mathbb{R}^{2n}$, writing $X=(x,\xi)\in\mathbb{R}^{2n}$ for the points in $\mathbb{R}^{2n}$. We write for short $\dot{\mathbb{R}}^{2n}=\mathbb{R}^{2n}\setminus\{0\}$. As usual, we put $\langle X\rangle=(1+|X|^{2})^{1/2}$, $D_{x_{j}}=-i\partial_{x_{j}}$. The $L^{2}$-inner product in $L^{2}(\mathbb{R}^{n})$ is denoted by $(\cdot,\cdot)_{0}$.
We next recall the Shubin calculus of global operator (in the context of the Weyl-Hörmander calculus).
@@
Let $a\in C^{\infty}(\mathbb{R}^{2n})$ and $m\in\mathbb{R}$. We say that $a$ is a global symbol of order $m$, and we write $a\in S^{m}(\mathbb{R}^{n})$, if for all $\alpha\in\mathbb{Z}_{+}^{2n}$ there exists $C_{\alpha}>0$
@@
One has that $a^{\mathrm{w}}(x,D)$ is a linear operator that acts continuously on $\mathscr{S}(\mathbb{R}^{n})$ and extends by duality to a linear continuous operator on $\mathscr{S}^{\prime}(\mathbb{R}^{n})$.
@@
We call the elements of $\Psi^{-\infty}(\mathbb{R}^{n})$ smoothing operators: they are those operators in $\Psi^{\infty}(\mathbb{R}^{n})$ that have smoothing - i.e. $\mathscr{S}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ - Schwartz kernel.
The formal adjoint of $a^{\mathrm{w}}(x,D)$ is simply $(\bar{a})^{\mathrm{w}}(x,D)$ and the composition law is written as follows: if $a\in S^{m}(\mathbb{R}^{n}$ and $b\in S^{m^{\prime}}(\mathbb{R}^{n})$ then $a^{\mathrm{w}}(x,D)b^{\mathrm{w}}(x,D)=(a\sharp b)^{\mathrm{w}}(x,D)$ where $a\sharp b\in S^{m+m^{\prime}}(\mathbb{R}^{n})$ is written, for any given $N\in\mathbb{Z}_{+}$, as
@@
We recall also the class $S^{m}_{\mathrm{sreg}}(\mathbb{R}^{n})$ of semiregular symbols.
@@
where $\mathsf{M}_{N}$ denotes the vector space of the $N\times N$ complex matrices. The composition formulas are then to be understood as compositions rows-times-columns, and the adjoint of a symbol will be given by the adjoint matrix of symbolic entries. Ellipticity is defined by requiring that $\det a\in S^{mN}(\mathbb{R}^{n})$ is elliptic. Analogously, one has the class of matrix-valued semiregular symbols $S^{m}_{\mathrm{sreg}}(\mathbb{R}^{n};\mathsf{M}_{N})$. In this case, ellipticity is a condition determined only by the principal symbol.
@@
As regards the $1$-d QRM we have the following setup. Given $p_{2}(X)=|X|^{2}/2$, $X\in\mathbb{R}^{2}$, we therefore have that $P_{0}=p_{2}^{\mathrm{w}}(x,D)\in\Psi^{2}(\mathbb{R})$ is the usual quantum harmonic oscillator $(D_{x}^{2}+x^{2})/2,$ $x\in\mathbb{R}.$ We have that $P_{0}$, realized as an unbounded operator in $L^{2}(\mathbb{R})$ with maximal domain the Shubin Sobolev space $B^{2}(\mathbb{R})$ (see [6, 24]), is self-adjoint and has a discrete spectrum of eigenvalues of multiplicity $1$ given by $\{N+1/2\}_{N\in\mathbb{Z}_{+}}$, and corresponding eigenfunctions, the ones we will use in our subsequent computations, given by the Schwartz functions
@@
Recall also, in passing, an important property of the Hermite polynomials:
@@
There exists a metaplectic operator $U\colon L^{2}(\mathbb{R};\mathbb{C}^{2})\longrightarrow L^{2}(\mathbb{R};\mathbb{C}^{2})$ (which is an automorphism of $\mathscr{S}(\mathbb{R};\mathbb{C}^{2})$ and $\mathscr{S}^{\prime}(\mathbb{R};\mathbb{C}^{2})$) such that
@@
3.1. Relation between $Q_{\mathrm{Rabi}}$ and $Q_{R}$
@@
$Q_{\mathrm{Rabi}}=p_{2}^{\mathrm{w}}(x,D)I_{2}+\alpha x\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]+\varepsilon\Delta\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right]-\frac{1}{2}I_{2},$
@@
$Q_{R}=Q_{\mathrm{Rabi}}+\frac{1}{2}I_{2},$
with $\gamma_{1}=-\gamma_{2}$ and $\gamma_{1}=\Delta$. Therefore up to translating the whole spectrum of $Q_{R}$ by $1/2$, studying the spectrum of $Q_{R}$ with $\Delta=\gamma_{1}=-\gamma_{2}$ is the same as studying the spectrum of $Q_{\mathrm{Rabi}}$.
@@
$U^{*}Q_{\mathrm{Rabi}}U=p_{2}^{\mathrm{w}}(x,D)I_{2}+\varepsilon\Delta\left[\begin{array}[]{cc}0&T_{+}^{2}\\ T_{-}^{2}&0\end{array}\right]-\frac{1+\alpha^{2}}{2}I_{2}.$
@@ 4. Perturbation Theory
because of the $L^{2}\to L^{2}$ continuity of $BE$ and $BR$. Therefore: The family $\varepsilon\mapsto A+\varepsilon B$ is a holomorphic regular family in the sense of Rellich [21] (equivalently, of type (A) in the sense of Kato) for all $|\varepsilon|\leq 1/(2C_{B})=\varepsilon_{0},$ see [9]. We denote by $I_{\varepsilon_{0}}$ the corresponding interval centered at 0. It follows that, by the very nature of $A+\varepsilon B$, it has a compact resolvent for all $\varepsilon$ and hence its eigenvalues can be described as analytic functions of $\varepsilon$ (see Kato [9], Thm. 3.9 page 392–393). We may therefore study the analyticity of finite families of eigenvalules of $Q_{R,\varepsilon}$ in $\varepsilon$.
@@
Consider the quadratic form $(B\cdot,\cdot)_{0}\bigl|_{E_{0}\times E_{0}}$: for linear combinations $u=w_{1}\boldsymbol{\phi}_{1}+w_{2}\boldsymbol{\phi}_{2}\in E_{0}$, $w=\left[\begin{array}[]{c}w_{1}\\ w_{2}\end{array}\right]\in\mathbb{C}^{2}$, and likewise for $u^{\prime}\in E_{0}$ with coefficients $w^{\prime}=\left[\begin{array}[]{c}w^{\prime}_{1}\\ w^{\prime}_{2}\end{array}\right]$, we have

### v2

@@ Abstract.
In the first part of the paper we study a perturbative model of the Rabi system of Quantum Optics. We are therefore able to describe, through Rellich’s theory, an analytic expansion of finite families of eigenvalues, of arbitrary fixed length. In particular, we prove that for finite families of eigenvalues the Braak conjecture holds. In the second part we study the asymptotics of the Weyl spectral counting function of a class of systems that generalize the Quantum Rabi Model to an $N$-level atom ($N\geq 3$) with N1N-1 cavity modes of the electromagnetic field.
@@ 1. Introduction
Being a system on $\mathbb{R}^{n}$ with polynomial coefficients, based on the Quantum Harmonic Oscillator, it fits in the framework of Non-Commutative Harmonic Oscillators (NCHO) introduced by Wakayama and Parmeggiani in [18, 19] and as later extended in Parmeggiani [16, 17], and more recently in Malagutti and Parmeggiani [13] (and [14]). However, a difficulty in the geometric analysis of the Rabi system is that it is not a semiregular globally elliptic system in the sense of [13], in that, although the symbol is principally scalar, the symbol of the first order part (the semiprincipal symbol) cannot be smoothly diagonalized in the phase space $\mathbb{R}^{n}\times\mathbb{R}^{n}$ (as instead is the case for the Jaynes-Cummings system).
@@
The paper is divided into two parts, both of which are perturbative in nature. In the first part we study a perturbation of the classical $1$-d QRM, which can be dealt with by the Rellich perturbation theory as in [22] (we chose that approach in place of Kato’s one [9] because more direct). In this part we destroy the homogeneity of the system, which allows us to bring the system into an isometric system to which we may apply perturbation theory. Since all the eigenvalues of the unperturbed system (a scalar quantum harmonic oscillator acting on $L^{2}(\mathbb{R};\mathbb{C}^{2})$) are degenerate, we apply Rellich’s approach by assuming some first-order (in the perturbation parameter) nondegeneracy, described by the way the interaction parameter is approximated by zeros of Legendre polynomials. That allows to control finite segments of the spectrum of the system, for a sufficiently small perturbation parameter, and to prove that in this context Braak’s Conjecture for $Q_{\mathrm{Rabi},\varepsilon}$, a particular case of our system $Q_{R}$ (see Section 5) holds, at least for arbitrarily long but finite parts of the spectrum. In addition, Rellich’s method gives control on the corresponding eigenfunctions. We also deal with ”degenerate” (in the sense of Rellich) situations, but in that case we use the simpler method of quasimodes (see Lazutkin [11]) that, however, gives up control on the eigenfunctions but retains control on the spectrum, proving its simplicity everywhere. Since in general we do not control in this case the eigenfunctions, we do not have information on the parity, which is basic for Braak’s conjecture. In all cases, our method, depending on the nondegeneracy of $Q^{(2)}_{N}$ (see Section 6), gives that by switching on the perturbation parameter $\varepsilon$ the perturbed eigenvalues may move to the right, or to the left, or both sides of the unperturbed eigenvalue $N$ (at $\varepsilon=0$) according to the signs of the eigenvalues of $Q^{(2)}_{N}$. This gives instances in which Braak’s conjecture keeps holding, however leaving open the possibility that in these degenerate cases Braak’s might no longer be true for very particular values of the interaction parameter. It also indicates a strategy to cope with more degenerate situations. In a sense, our result in the nondegenerate case may be thought of as a complement of Rudnik’s result [23]. For generalizations of the Braak’s Conjecture to asymmetric quantum Rabi models, see the recent paper by Braak, Nguyen, Reyes-Bustos and Wakayama [3].
@@
In the second part of the paper we describe in the first place some possible generalizations of the QRM to higher dimensions for an interaction with atoms having several energy levels, where the space dimension $n$ is related to the number $N$ of atomic levels by the equation n=N1.n=N-1. In order to approach the study of the asymptotics of the Weyl spectral counting function, we have to cope with the problem of the non-diagonalizability in the symbol class of the first order part of the system. We then use a perturbation method that brings the perturbed symbol into the SMGES class, for which we may exploit a pseudodifferential diagonalization of the system and the results of [13]. Variational considerations then allow us to give a two-term asymptotics of the Weyl function. The result is not known in the higher dimensional and energy level number case.
@@
In Section 2 we describe the Quantum Rabi Model from a physical perspective, highlighting its connection with the Jaynes-Cummings model by Rotating Wave Approximation and its importance as a foundational model in Quantum Optics. In Section 3 we recall our setup, and review the global pseudodifferential calculus used especially in the final Section 7. In Section 4 we give the perturbative result, under a non-degeneration condition, on finite initial segments of the spectrum of the QRM. In Section 5 we recall Braak’s conjecture and show its validity in our setting. We next deal with the degenerate case by the use of quasi-modes, which is the content of Section 6. We will show instances in which Braak’s conjecture keeps holding also in the degenerate case. In the final Section 7 we introduce the generalizations of the QRM and give a perturbative proof of a two-term asymptotics of the Weyl spectral counting function.
@@ 2. Physics of the Quantum Rabi Model
The Quantum Rabi Model (QRM) describes the fundamental interaction between a two-level quantum system (a qubit, atom, or spin-½) and a single bosonic mode (a quantized harmonic oscillator, typically a cavity or vibrational mode). It plays a central role in Quantum Optics, cavity and circuit QED, trapped-ion system, and modern quantum technologies, as it captures the minimal setting in which light-matter interaction is in full treated quantum mechanically.
@@ 2.2. Symmetry and integrability
which commutes with the Hamiltonian. This symmetry splits the Hilbert space into even and odd parity subspaces and is essential for understanding the structure of the spectrum. Braak [1] showed that the QRM is integrable in a generalized sense and that its spectrum can be obtained from the zeros of transcendental functions having no simple closed-form expression, but built from infinite series and special functions.
@@ 2.3.1. Weak coupling: Jaynes-Cummings regime
For $g\ll\omega,\Omega$, and near resonance $\omega\approx\Omega$, the rapidly oscillating counter-rotating terms can be neglected under the rotating-wave approximation (RWA). The Hamiltonian reduces to
@@ 2.3.2. Ultrastrong coupling
When $g/\omega=0.1\pm\epsilon$, the counter-rotating terms become relevant and the RWA breaks down. This is the ultrastrong coupling regime, experimentally realized in superconducting circuit QED (see the paper by Niemczyk [15]). The ground state contains virtual photons and the system exhibits phenomena such as the Bloch–Siegert shift.
@@ 3. The setup
We work in the phase-space $T^{\ast}\mathbb{R}^{n}=\mathbb{R}^{n}\times\mathbb{R}^{n}=\mathbb{R}^{2n}$, writing $X=(x,\xi)\in\mathbb{R}^{2n}$ for the points in $\mathbb{R}^{2n}$. We write for short $\dot{\mathbb{R}}^{2n}=\mathbb{R}^{2n}\setminus\{0\}$. As usual, we put $\langle X\rangle=(1+|X|^{2})^{1/2}$, $D_{x_{j}}=-i\partial_{x_{j}}$. The $L^{2}$-inner product in $L^{2}(\mathbb{R}^{n})$ (or $L^{2}(\mathbb{R}^{n};\mathbb{C}^{N})$) is denoted by $(\cdot,\cdot)_{0}$.
We next recall the Shubin calculus of global operators (in the context of the Weyl-Hörmander calculus).
@@
Let $a\in C^{\infty}(\mathbb{R}^{2n})$ and $m\in\mathbb{R}$. We say that $a$ is a global symbol of order $m$, and write $a\in S^{m}(\mathbb{R}^{n})$, if for all $\alpha\in\mathbb{Z}_{+}^{2n}$ there exists $C_{\alpha}>0$ such that
@@
One has that $a^{\mathrm{w}}(x,D)$ is a linear operator that acts continuously $\mathscr{S}(\mathbb{R}^{n})\to\mathscr{S}(\mathbb{R}^{n})$ and extends by duality to a linear continuous operator $\mathscr{S}^{\prime}(\mathbb{R}^{n})\to\mathscr{S}^{\prime}(\mathbb{R}^{n})$.
@@
We call the elements of $\Psi^{-\infty}(\mathbb{R}^{n})$ smoothing operators: they are those operators in $\Psi^{\infty}(\mathbb{R}^{n})$ that have smoothing, i.e. $\mathscr{S}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, Schwartz kernel.
The formal adjoint of $a^{\mathrm{w}}(x,D)$ is simply ${\bar{a}}^{\mathrm{w}}(x,D)$ and the composition law is written as follows: If $a\in S^{m}(\mathbb{R}^{n}$ and $b\in S^{m^{\prime}}(\mathbb{R}^{n})$ then $a^{\mathrm{w}}(x,D)b^{\mathrm{w}}(x,D)=(a\sharp b)^{\mathrm{w}}(x,D)$ where $a\sharp b\in S^{m+m^{\prime}}(\mathbb{R}^{n})$ is written, for any given $N\in\mathbb{Z}_{+}$, as
@@
We recall also the important class $S^{m}_{\mathrm{sreg}}(\mathbb{R}^{n})$ of semiregular symbols.
@@
where $\mathsf{M}_{N}$ denotes the space of the $N\times N$ complex matrices. The composition formulas are then to be understood as compositions rows-times-columns, and the adjoint of a symbol will be given by the adjoint matrix of symbolic entries. Ellipticity is defined by requiring that $\det a\in S^{mN}(\mathbb{R}^{n})$ is elliptic. Analogously, one has the class of matrix-valued semiregular symbols $S^{m}_{\mathrm{sreg}}(\mathbb{R}^{n};\mathsf{M}_{N})$. In this case, ellipticity is a condition determined only by the principal symbol.
@@
Next, as regards the $1$-d QRM we have the following setup. Given $p_{2}(X)=|X|^{2}/2$, $X\in\mathbb{R}^{2}$, we have that $P_{0}=p_{2}^{\mathrm{w}}(x,D)\in\Psi^{2}(\mathbb{R})$ is the usual quantum harmonic oscillator $(D_{x}^{2}+x^{2})/2,$ $x\in\mathbb{R}.$ We have that $P_{0}$, realized as an unbounded operator in $L^{2}(\mathbb{R})$ with maximal domain the Shubin Sobolev space $B^{2}(\mathbb{R})$ (see [6, 25]), is self-adjoint and has a discrete spectrum made of eigenvalues with multiplicity $1$ given by the sequence $\{N+1/2\}_{N\in\mathbb{Z}_{+}}$, and corresponding eigenfunctions, the ones we will use in our subsequent computations, given by the Schwartz functions
@@
Recall also, in passing, an important property of the Hermite polynomials that will be needed shortly:
@@
There exists a metaplectic operator $U\colon L^{2}(\mathbb{R};\mathbb{C}^{2})\longrightarrow L^{2}(\mathbb{R};\mathbb{C}^{2})$ (which is an automorphism of $\mathscr{S}(\mathbb{R};\mathbb{C}^{2})$ and of $\mathscr{S}^{\prime}(\mathbb{R};\mathbb{C}^{2})$) such that
@@
3.1. Relation between $Q_{\mathrm{Rabi},\varepsilon}$ and $Q_{R}$
@@
$Q_{\mathrm{Rabi},\varepsilon}=p_{2}^{\mathrm{w}}(x,D)I_{2}+\alpha x\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]+\varepsilon\Delta\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right]-\frac{1}{2}I_{2},$
@@
$Q_{R,\varepsilon}=Q_{\mathrm{Rabi},\varepsilon}+\frac{1}{2}I_{2},$
with $\gamma_{1}=-\gamma_{2}$ and $\gamma_{1}=\Delta$. Therefore up to translating the whole spectrum of $Q_{R}$ by $1/2$, studying the spectrum of $Q_{R,\varepsilon}$ with $\Delta=\gamma_{1}=-\gamma_{2}$ is the same as studying the spectrum of $Q_{\mathrm{Rabi},\varepsilon}$.
@@
$U^{*}Q_{\mathrm{Rabi},\varepsilon}U=p_{2}^{\mathrm{w}}(x,D)I_{2}+\varepsilon\Delta\left[\begin{array}[]{cc}0&T_{+}^{2}\\ T_{-}^{2}&0\end{array}\right]-\frac{1+\alpha^{2}}{2}I_{2}.$
@@ 4. Perturbation Theory
because of the $L^{2}\to L^{2}$ continuity of $BE$ and $BR$. Therefore: The family $\varepsilon\mapsto A+\varepsilon B$ is a holomorphic regular family in the sense of Rellich [22] (equivalently, of type (A) in the sense of Kato) for all $|\varepsilon|\leq 1/(2C_{B})=:\varepsilon_{0},$ see [9]. We denote by $I_{\varepsilon_{0}}$ the corresponding interval centered at 0. It follows that, by the very nature of $A+\varepsilon B$, it has a compact resolvent for all $\varepsilon$ and hence its eigenvalues can be described as analytic functions of $\varepsilon$ (see Kato [9], Thm. 3.9 page 392–393). We may therefore study the analyticity of finite families of eigenvalues of $Q_{R,\varepsilon}$ in $\varepsilon$.
@@
(Therefore $R_{0}$ is a parametrix of the globally elliptic operator $A-\lambda_{0},$ because $\pi_{E_{0}}$ is a smoothing term since it has rapidly decreasing Schwartz kernel.)
Consider the bilinear form $(B\cdot,\cdot)_{0}\bigl|_{E_{0}\times E_{0}}$: for linear combinations $u=w_{1}\boldsymbol{\phi}_{1}+w_{2}\boldsymbol{\phi}_{2}\in E_{0}$, $w=\left[\begin{array}[]{c}w_{1}\\ w_{2}\end{array}\right]\in\mathbb{C}^{2}$, and likewise for $u^{\prime}\in E_{0}$ with coefficients $w^{\prime}=\left[\begin{array}[]{c}w^{\prime}_{1}\\ w^{\prime}_{2}\end{array}\right]$, we have